OpenStudy (anonymous):

Please help Which of the following is the vertex of the quadratic equation y = –4(x + 6)2 + 2

OpenStudy (anonymous):

the vertex of a quadratic written in this form: \(\large y=a(x-h)^2+k \) is (h, k)

OpenStudy (anonymous):

huhh?

OpenStudy (anonymous):

just match it up with your equation...

OpenStudy (anonymous):

h = x and k=y right?

OpenStudy (anonymous):

just match up the equation with your equation... what matches the h, and k ? \[\huge y=a(x-h)^2+k \] \[\huge y = –4(x + 6)^2 + 2 \]

OpenStudy (anonymous):

ok..im stupid at math, so can you walk me through the steps

OpenStudy (anonymous):